# HyperMono: A Monotonicity-aware Approach to Hyper-Relational Knowledge Representation

### v1

@@ 4.2 MEP: Missing Entity Predictor
    $\left\{\begin{array}[]{l}\alpha_{h}=f(\texttt{MLP}(\textbf{{E}}^{h})),\,\,\beta_{h}=g(\texttt{MLP}(\textbf{{E}}^{h}))\vspace{1.2ex}\\ \alpha_{r}=f(\texttt{MLP}(\textbf{{E}}^{r})),\,\,\beta_{r}=g(\texttt{MLP}(\textbf{{E}}^{r}))\vspace{1.2ex}\\ f(x_{i})=\pi\texttt{tanh}(\lambda_{1}x_{i}),\,\,g(y_{i})=\pi\texttt{tanh}(\lambda_{2}y_{i})+\pi\\ \end{array}\right.$ (2)
@@
    $\left\{\begin{array}[]{l}\alpha_{r,a_{i},v_{i}}=f(\texttt{MLP}_{1}(\Theta_{r,a_{i},v_{i}}))\vspace{1.2ex}\\ \beta_{r,a_{i},v_{i}}=g(\texttt{MLP}_{1}(\Theta_{r,a_{i},v_{i}}))\vspace{1.2ex}\\ {\rm offset}=(\beta_{h,r}-\beta_{\mathcal{S}_{i}})\odot\sigma(\alpha_{r,a_{i},v_{i}}),\,\,\beta_{\mathcal{S}_{i}}=\sigma(\beta_{r,a_{i},v_{i}})\vspace{1.2ex}\\ \alpha_{\mathcal{S}_{i}}=\alpha_{h,r}-0.5\odot\beta_{h,r}+0.5\odot\beta_{\mathcal{S}_{i}}+0.5\odot{\rm offset}\\ \end{array}\right.$ (3)
@@
    Figure 3: An overview of cone embedding with projection $\mathcal{P}$, intersection $\mathcal{I}$ and shrink $\mathcal{S}$, the embedding dimension $d$=1.
@@ 4.3 Joint Training
$\mathcal{L}_{\textit{joint}}=\mathcal{L}_{\textit{triple}}(h)+\mathcal{L}_{\textit{hyper}}(h)+\mathcal{L}_{\textit{triple}}(t)+\mathcal{L}_{\textit{hyper}}(t)$ (4)
5 Experiments
@@
  • RQ3 (Model transferability): Can HyperMono be used to complete hypergraphs?
  • RQ4 (Parameter sensitivity): How hyper-parameters influence its performance?
@@
  • RQ5 (Complexity analysis): What is the amount of computation and parameters used by HyperMono?
5.1 Experiment Setup
5.1.1 Datasets
@@
We conduct experiments on three hyper-relational KGs: WD50K [17], WikiPeople [41], and JF17K [13], with WD50K and WikiPeople being extracted from Wikidata [42] and JF17K from Freebase [43]. These datasets have the following two characteristics. On the one hand, not all main triples contain hyper-relational knowledge: only 13.6% in WD50K, 2.6% in WikiPeople and 45.9% in JF17K. On the other hand, each main triple has a different number of qualifier pairs: ranging from 0 to 20 in WD50K, 0 to 7 in WikiPeople and 0 to 4 in JF17K. To validate the models’ robustness across multiple scenarios, these datasets have been refined taking into account the two described characteristics [23], i.e., the percentage of triples containing hyper-relational knowledge (mixed-percentage or fixed-percentage) and the number of qualifiers associated to main triples (mixed-qualifier or fixed-qualifier). We can thus construct three different sub-datasets, namely: Mixed-percentage Mixed-qualifier, Fixed-percentage Mixed-qualifier, and Fixed-percentage Fixed-qualifier. Note that the by definition Mixed-percentage Fixed-qualifier scenario does not exist. The statistics of the considered datasets are as shown in the Table I.
  1. 1.
    Mixed-percentage Mixed-qualifier. This is the original setting introduced in [17], used to test the performance under variable number of main triples with qualifiers and with different number of qualifiers associated to those triples.
  2. 2.
    Fixed-percentage Mixed-qualifier. Following [17, 23], we construct subsets where the number of qualifiers for each triple varies, but the proportion of triples with hyper-relational knowledge is fixed. For example, in the subsets WD50K (33), WD50K (66), and WD50K (100), the number in parentheses indicates the corresponding proportion of main triples with hyper-relational knowledge. We consider similar subsets for the WikiPeople and JF17K datasets.
  3. 3.
@@
TABLE I: Statistics of datasets under three different scenarios. The values in parentheses indicate the percentage of triples in the corresponding dataset with hyper-relational facts. The hyphen -n indicates that each main triple in the corresponding dataset contains n qualifier pairs.
@@
WikiPeople-3 20656 2582 2582 12270 66
WikiPeople-4 12150 1519 1519 9528 50
JF17K-3 27635 3454 3455 11541 104
JF17K-4 7607 951 951 6536 23
TABLE II: Evaluation of different models under the mixed-percentage mixed-qualifier scenario on the WD50K, WikiPeople and JF17K datasets. All baseline results are collected from the original papers. Best scores are highlighted in bold, the second best scores are underlined, and ’–’ indicates the results are not reported in previous work. The number in parentheses represents the proportion of triples with hyper-relational knowledge.
Methods WD50K (13.6) WikiPeople (2.6) JF17K (45.9)
MRR Hits@1 Hits@10 MRR Hits@1 Hits@10 MRR Hits@1 Hits@10
Embedding-based Methods
m-TransH [13] 0.063 0.063 0.300 0.206 0.206 0.463
RAE [14] 0.059 0.059 0.306 0.215 0.215 0.469
NaLP-Fix [15] 0.177 0.131 0.264 0.420 0.343 0.556 0.245 0.185 0.358
HINGE [15] 0.243 0.176 0.377 0.476 0.415 0.585 0.449 0.361 0.624
ShrinkE [16] 0.345 0.275 0.482 0.485 0.431 0.601 0.589 0.506 0.749
Transformer-based with GNNs Methods
StarE [17] 0.349 0.271 0.496 0.491 0.398 0.648 0.574 0.496 0.725
QUAD [18] 0.348 0.270 0.497 0.466 0.365 0.624 0.582 0.502 0.740
QUAD (Parallel) [18] 0.349 0.275 0.489 0.497 0.431 0.617 0.596 0.519 0.751
Transformer-based without GNNs Methods
HyTrans [19] 0.356 0.281 0.498 0.501 0.426 0.634 0.582 0.501 0.742
GRAN-hete [20] 0.503 0.438 0.620 0.617 0.539 0.770
GRAN-homo [20] 0.487 0.410 0.618 0.611 0.533 0.767
GRAN-complete [20] 0.489 0.413 0.617 0.591 0.510 0.753
HyNT [21] 0.314 0.241 0.454 0.464 0.380 0.601 0.554 0.475 0.711
HAHE [22] 0.368 0.291 0.516 0.509 0.447 0.639 0.623 0.554 0.806
HyperFormer [23] 0.366 0.288 0.514 0.491 0.386 0.654 0.664 0.601 0.787
HyperMono 0.375 0.298 0.522 0.494 0.390 0.657 0.685 0.630 0.796
5.1.2 Baselines, Evaluation Protocol and Implementation Details
Baselines. We compare HyperMono with twelve state-of-the-art baselines for HKGC, including Embedding-based Methods, Transformer-based with GNNs Methods, and Transformer-based without GNNs Methods:

### v2

@@ 3.2.2. MEP: Missing Entity Predictor
Figure 3. An overview of cone embedding with projection $\mathcal{P}$, intersection $\mathcal{I}$ and shrink $\mathcal{S}$, the embedding dimension $d$=1.
@@
    (2) $\left\{\begin{array}[]{l}\alpha_{h}=f(\texttt{MLP}(\textbf{{E}}^{h})),\,\,\beta_{h}=g(\texttt{MLP}(\textbf{{E}}^{h}))\vspace{1.2ex}\\ \alpha_{r}=f(\texttt{MLP}(\textbf{{E}}^{r})),\,\,\beta_{r}=g(\texttt{MLP}(\textbf{{E}}^{r}))\vspace{1.2ex}\\ f(x_{i})=\pi\texttt{tanh}(\lambda_{1}x_{i}),\,\,g(y_{i})=\pi\texttt{tanh}(\lambda_{2}y_{i})+\pi\\ \end{array}\right.$
@@
    (3) $\left\{\begin{array}[]{l}\alpha_{r,a_{i},v_{i}}=f(\texttt{MLP}_{1}(\Theta_{r,a_{i},v_{i}})),\,\,\beta_{r,a_{i},v_{i}}=g(\texttt{MLP}_{1}(\Theta_{r,a_{i},v_{i}}))\vspace{1.2ex}\\ {\rm offset}=(\beta_{h,r}-\beta_{\mathcal{S}_{i}})\odot\sigma(\alpha_{r,a_{i},v_{i}}),\,\,\beta_{\mathcal{S}_{i}}=\sigma(\beta_{r,a_{i},v_{i}})\vspace{1.2ex}\\ \alpha_{\mathcal{S}_{i}}=\alpha_{h,r}-0.5\odot\beta_{h,r}+0.5\odot\beta_{\mathcal{S}_{i}}+0.5\odot{\rm offset}\\ \end{array}\right.$
@@ 3.2.3. Joint Training
(4) $\mathcal{L}_{\textit{joint}}=\mathcal{L}_{\textit{triple}}(h)+\mathcal{L}_{\textit{hyper}}(h)+\mathcal{L}_{\textit{triple}}(t)+\mathcal{L}_{\textit{hyper}}(t)$
4. Experiments
@@
  • RQ3 (Parameter sensitivity): How hyper-parameters influence its performance? We present part of the experimental results in the main body, further detailed analysis can be found in Appendix D.
@@
  • RQ4 (Complexity analysis): What is the amount of computation and parameters used by HyperMono?
  • RQ5 (Model transferability): Can HyperMono be used to complete hypergraphs? This results can be found in Appendix E.
Table 1. Evaluation of different models under the fixed-percentage fixed-qualifier scenario on the WikiPeople and JF17K datasets. Best scores are highlighted in bold, the second best scores are highlighted in normal.
Methods WikiPeople-3 WikiPeople-4 JF17K-3 JF17K-4
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
ShrinkE (Xiong et al., 2023) 0.366 0.291 0.397 0.515 0.222 0.155 0.252 0.353 0.712 0.648 0.745 0.837 0.779 0.726 0.809 0.886
StarE (Galkin et al., 2020) 0.401 0.310 0.434 0.592 0.243 0.156 0.269 0.430 0.707 0.635 0.744 0.847 0.723 0.669 0.753 0.839
QUAD (Wang et al., 2021) 0.403 0.321 0.438 0.563 0.251 0.167 0.280 0.425 0.730 0.660 0.767 0.870 0.787 0.730 0.823 0.895
Hy-Transformer (Yu and Yang, 2021) 0.403 0.323 0.436 0.569 0.248 0.165 0.275 0.422 0.690 0.617 0.725 0.837 0.773 0.717 0.806 0.875
GRAN (Wang et al., 2021) 0.397 0.328 0.429 0.533 0.239 0.178 0.261 0.364 0.779 0.724 0.811 0.893 0.798 0.744 0.830 0.904
HyNT (Chung et al., 2023) 0.383 0.306 0.415 0.533 0.233 0.166 0.257 0.369 0.733 0.664 0.771 0.873 0.779 0.725 0.809 0.881
HAHE (Luo et al., 2023) 0.397 0.323 0.432 0.537 0.251 0.179 0.286 0.392 0.762 0.701 0.791 0.892 0.736 0.666 0.773 0.869
HyperFormer (Hu et al., 2023) 0.573 0.511 0.603 0.693 0.393 0.336 0.415 0.496 0.832 0.790 0.855 0.914 0.857 0.811 0.884 0.937
4.1. Experiment Setup
4.1.1. Datasets
We conduct experiments on three hyper-relational KGs: WD50K (Galkin et al., 2020), WikiPeople (Guan et al., 2019), and JF17K (Wen et al., 2016), where WD50K and WikiPeople have been extracted from Wikidata (Vrandecic and Krötzsch, 2014) and JF17K from Freebase (Bollacker et al., 2008). We also construct three different sub-datasets, namely: Fixed-percentage Fixed-qualifier, Fixed-percentage Mixed-qualifier, and Mixed-percentage Mixed-qualifier. Statistics and details on the construction can be found in Appendix A.
4.1.2. Baselines.
We compare HyperMono with twelve state-of-the-art baselines for HKGC, including Embedding-based Methods: m-TransH (Wen et al., 2016), RAE (Zhang et al., 2018), NaLP-Fix (Rosso et al., 2020), HINGE (Rosso et al., 2020), and ShrinkE (Xiong et al., 2023); Transformer-based with GNNs Methods: StarE (Galkin et al., 2020) and QUAD (Shomer et al., 2022); Transformer-based without GNNs Methods: Hy-Transformer (Yu and Yang, 2021), GRAN (Wang et al., 2021), HyNT (Chung et al., 2023), HAHE (Luo et al., 2023), and HyperFormer (Hu et al., 2023). Detailed baseline descriptions, evaluation metrics and implementation details can respectively be found in Appendix B and C.
Table 2. Evaluation of different models under the fixed-percentage mixed-qualifier scenario on the WD50K, WikiPeople and JF17K datasets. Best scores are highlighted in bold, the second best scores are highlighted in normal.
Methods WD50K WikiPeople JF17K
33% 66% 100% 33% 66% 100% 33% 66% 100%
MRR H@1 MRR H@1 MRR H@1 MRR H@1 MRR H@1 MRR H@1 MRR H@1 MRR H@1 MRR H@1
ShrinkE (Xiong et al., 2023) 0.246 0.190 0.356 0.294 0.470 0.404 0.176 0.129 0.244 0.198 0.332 0.276 0.259 0.176 0.263 0.185 0.299 0.220
StarE (Galkin et al., 2020) 0.308 0.247 0.449 0.388 0.610 0.543 0.192 0.143 0.259 0.205 0.343 0.279 0.290 0.197 0.302 0.214 0.321 0.223
QUAD (Wang et al., 2021) 0.329 0.266 0.479 0.416 0.646 0.572 0.204 0.155 0.282 0.228 0.385 0.318 0.307 0.210 0.334 0.241 0.379 0.277
Hy-Transformer (Yu and Yang, 2021) 0.313 0.255 0.458 0.397 0.621 0.557 0.192 0.140 0.268 0.215 0.372 0.316 0.298 0.204 0.325 0.234 0.361 0.266
GRAN (Wang et al., 2021) 0.322 0.269 0.472 0.419 0.647 0.593 0.201 0.156 0.287 0.244 0.403 0.349 0.307 0.212 0.326 0.237 0.382 0.290
HyNT (Chung et al., 2023) 0.295 0.238 0.438 0.377 0.595 0.525 0.175 0.130 0.243 0.192 0.329 0.269 0.285 0.195 0.298 0.216 0.352 0.261
HAHE (Luo et al., 2023) 0.301 0.239 0.432 0.368 0.607 0.537 0.188 0.142 0.263 0.213 0.367 0.304 0.311 0.221 0.330 0.248 0.387 0.293
HyperFormer (Hu et al., 2023) 0.338 0.280 0.492 0.434 0.666 0.611 0.213 0.161 0.298 0.255 0.426 0.373 0.352 0.254 0.411 0.325 0.478 0.396
4.2. Main Results
To address RQ1, we conduct experiments on the three types of datasets described above. The corresponding results are shown in Tables 3, 2, and 1.
Fixed-percentage Fixed-qualifier. To better understand the impact of hyper-knowledge on the performance, we consider the WikiPeople and JF17K datasets with a fixed number of qualifiers, cf. Table 1. We can observe that HyperMono consistently achieves optimal performance on almost all datasets. Specifically, in terms of the MRR metric, HyperMono respectively shows improvements of 3.5% and 2.4% on the JF17K-3 and JF17K-4 subsets when compared with HyperFormer. Furthermore, we find that for all models, their performance in the fixed-percentage fixed-qualifier setting is generally higher than in the fixed-percentage mixed-qualifier one. The main reason is that in the fixed-percentage fixed-qualifier scenario, the number of qualifier pairs for each triple is fixed, and there is no imbalance in the distribution of hyper-relational knowledge.
Fixed-percentage Mixed-qualifier. Table 2 presents the results on WD50K, WikiPeople, and JF17K under different fixed ratios of relational facts with qualifiers. For each dataset, we construct three subsets (as describe in § 4.1.1, Point 2) containing approximately 33%, 66%, and 100% of facts with qualifiers. We observe that under different proportions of hyper-relational knowledge HyperMono also achieves optimal performance. Particularly, for the WikiPeople dataset, in which HyperMono performs poorly in the mixed-percentage mixed-qualifier scenario, adding a certain percentage of qualifiers information to triples results in HyperMono outperforming all baselines. For example, for the Hits@1 metric, HyperMono respectively achieves improvements over the best performing baseline, HyperFormer, of 1.2% / 0.9% / 0.6% in the 33% / 66% / 100% variants. This confirms that the proportion of hyper-relational knowledge is a major factor influencing the performance over WikiPeople under the mixed-percentage mixed-qualifier condition. In contrast, on the WD50K dataset, no significant additional improvement was obtained. The most likely reason is the insufficient richness of hyper-relational knowledge in WD50K. For example, in WD50k_100 65.23% of the triples only contain one hyper-relation pair, which limits the functionality of qualifier monotonicity module.
Table 3. Evaluation of different models under the mixed-percentage mixed-qualifier scenario on the WD50K, WikiPeople and JF17K datasets. All baseline results are collected from the original papers. Best scores are highlighted in bold, the second best scores are highlighted in normal, and ’–’ indicates the results are not reported in previous work. The number in parentheses represents the proportion of triples with hyper-relational knowledge.
Methods WD50K (13.6) WikiPeople (2.6) JF17K (45.9)
MRR Hits@1 Hits@10 MRR Hits@1 Hits@10 MRR Hits@1 Hits@10
m-TransH (Wen et al., 2016) 0.063 0.063 0.300 0.206 0.206 0.463
RAE (Zhang et al., 2018) 0.059 0.059 0.306 0.215 0.215 0.469
NaLP-Fix (Rosso et al., 2020) 0.177 0.131 0.264 0.420 0.343 0.556 0.245 0.185 0.358
HINGE (Rosso et al., 2020) 0.243 0.176 0.377 0.476 0.415 0.585 0.449 0.361 0.624
ShrinkE (Xiong et al., 2023) 0.345 0.275 0.482 0.485 0.431 0.601 0.589 0.506 0.749
StarE (Galkin et al., 2020) 0.349 0.271 0.496 0.491 0.398 0.648 0.574 0.496 0.725
QUAD (Shomer et al., 2022) 0.349 0.275 0.489 0.497 0.431 0.617 0.596 0.519 0.751
Hy-Transformer (Yu and Yang, 2021) 0.356 0.281 0.498 0.501 0.426 0.634 0.582 0.501 0.742
GRAN (Wang et al., 2021) 0.489 0.413 0.617 0.591 0.510 0.753
HyNT (Chung et al., 2023) 0.314 0.241 0.454 0.464 0.380 0.601 0.554 0.475 0.711
HAHE (Luo et al., 2023) 0.368 0.291 0.516 0.509 0.447 0.639 0.623 0.554 0.806
HyperFormer (Hu et al., 2023) 0.366 0.288 0.514 0.491 0.386 0.654 0.664 0.601 0.787
HyperMono 0.375 0.298 0.522 0.494 0.390 0.657 0.685 0.630 0.796
Mixed-percentage Mixed-qualifier. Table 3 summarizes the results on WD50K, WikiPeople, and JF17K under the mixed-percentage mixed-qualifier condition. We can observe that HyperMono outperforms existing SoTA baselines by a large margin across most metrics on WD50K and JF17K. In particular, when compared with HyperFormer (the best performing baseline), HyperMono respectively achieves 2.1% and 2.9% improvements in MRR and Hits@1 on the JF17K dataset. A similar phenomenon occurred on the WD50K dataset. However, for the WikiPeople dataset, even though we can achieve optimal performance on the Hits@10 metric, we cannot surpass the best performing baseline on MRR and Hits@1. The intuitive reason for this is that, in comparison to the WD50K and JF17K datasets where 13.6% and 45.9% of main triple respectively contain hyper-relational knowledge, only 2.6% of triples in WikiPeople have qualifier pairs. This hinders HyperMono’s ability to exploit stage reasoning and qualifier monotonicity. Specifically, we employ a scaling transformation in the cone space to capture qualifier monotonicity, but the lack of qualifier pairs constrains this functionality.
Table 4. Ablation studies about different components. Best scores are highlighted in bold.
Methods WikiPeople-3 WikiPeople-4 JF17K-3 JF17K-4
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
w/o CNA+TP 0.559 0.493 0.591 0.686 0.367 0.303 0.382 0.477 0.832 0.792 0.855 0.912 0.860 0.820 0.885 0.940
w/o GEI 0.584 0.527 0.606 0.687 0.396 0.355 0.404 0.471 0.866 0.838 0.878 0.926 0.877 0.842 0.898 0.941
4.3. Ablation Studies
To address RQ2, we conduct ablation experiments on WikiPeople-3, WikiPeople-4, JF17K-3, and JF17K-4 subsets to verify the contribution of each component of HyperMono. These include the following three aspects: a) whether the knowledge of qualifiers is considered to infer missing entities, see rows “w/o CNA+TP” and “w/o FNA+QMP” in Table 4; b) different neighbor aggregation strategies, see rows “w/o LEI” and “w/o GEI” in Table 4; c) without the cone shrink operation in QMP, see row “w/o CSB” in Table 4.
Different Variants of HyperMono. In Table 4, we observe that under most metrics removing any functional module will bring some performance degradation, which justifies the necessity of each component. Specifically, a) removing either CNA+TP, which are related to coarse-grained answer prediction, or FNA+QMP, which are related to with fine-grained answer prediction, will result in a significant loss of accuracy. The intuitive reason is that the coarse-grained answer set can be used to constrain the decision space of fine-grained answers, preventing the occurrence of answer drift. Meanwhile, the fine-grained answer set allows to refine the answer set, ensuring the reliability of the predicted results. b) The neighborhood information of known entities has a positive effect on predicting missing entities, and the use of LEI for integrating the neighbors of known entities has a greater impact on inference than GEI. c) Removing the CSB module will bring a significant decrease in performance. The main reason is that the CSB module captures the qualifier monotonicity of hyper-relational KGs. The main characteristic of the HKGC task is the application of qualifier knowledge in triples to infer answer sets. Through the natural scaling transformation of the cone space, it can be captured that attaching qualifiers to a triple can only narrow down the answer set, but never enlarge it.
4.4. Parameter Sensitivity
To address RQ3, we present some of the parameter sensitivity experiments on WikiPeople-3, including: a) effect of label smoothing; b) effect of embedding dimension $d$. A more detailed analysis of other parameters (i.e., effect of temperature $\gamma$ and effect of initialization range) and other datasets (i.e., WikiPeople-4, JF17K-3, and JE17K-4 datasets) can be found in Appendix D. From Figure 4(a), we observe that the values of label smoothing have an impact on the performance, but there is no evident pattern. However, as shown in Figure 4(b), there is not a strict proportional relationship between embedding dimensions and the actual performance. After reaching a certain value of embedding dimensions, the accuracy stabilizes.
(a) Label Smoothing
(b) Embedding Dimension $d$
Figure 4. Parameter sensitivity experiments under different conditions on Wikipeople-3 dataset.
4.5. Complexity Analysis
To address RQ4, we analyze HyperMono’s complexity on the WD50K, WikiPeople, and JF17K datasets from two perspectives: time complexity and space complexity. The corresponding results are shown in Table 5. The time complexity is measured by the number of floating-point operations (#FLOPs) required during training phase, while the space complexity refers to the amount of a model’s parameters (#Params). The larger the values of #FLOPs and #Params are, the more computing power and higher memory usage are required during the training process. We observe that compared with HyperFormer, HyperMono requires less time and space overhead in each dataset. Specifically, on the WD50K dataset, HyperMono only requires nearly half the #FLOPs and one-sixth the #Params of HyperFormer, which demonstrates that HyperMono is more efficient and can achieve better results while occupying less time and space.
Table 5. Amount of parameters and calculation of the HyperFormer and HyperMono models.
Metrics Model WD50K WikiPeople JF17K
#FLOPs HyperFormer 118.397G 118.167G 118.070G
HyperMono 69.600G 69.371G 69.273G
#Params HyperFormer 66.956M 66.956M 66.956M
HyperMono 10.645M 10.645M 10.645M
5. Conclusion and Future Work
In this paper, we propose HyperMono, a framework for hyper-relational graph completion that properly implements two-stage reasoning and faithfully captures qualifier monotonicity. Specifically, we model each qualifier as a shrinking of the main triple cone to a qualifier cone to realize qualifier monotonicity. Empirical experiments under various settings on the WD50K, WikiPeople, and JF17K datasets demonstrate HyperMono’s robust performance, when compared to competitive baselines. Additionally, we present various ablation studies on parameter sensitivity, complexity analysis and model transferability. For future work, we believe that it is necessary to pay special attention to numerical attribute knowledge in hyper-relational knowledge. How to elegantly integrate schema knowledge into the representation process is also an interesting line of research.
A   Datasets
@@
We conduct experiments on three hyper-relational KGs: WD50K (Galkin et al., 2020), WikiPeople (Guan et al., 2019), and JF17K (Wen et al., 2016), with WD50K and WikiPeople being extracted from Wikidata (Vrandecic and Krötzsch, 2014) and JF17K from Freebase (Bollacker et al., 2008). These datasets have the following two characteristics. On the one hand, not all main triples contain hyper-relational knowledge: only 13.6% in WD50K, 2.6% in WikiPeople and 45.9% in JF17K. On the other hand, each main triple has a different number of qualifier pairs: ranging from 0 to 20 in WD50K, 0 to 7 in WikiPeople and 0 to 4 in JF17K. To validate the models’ robustness across multiple scenarios, these datasets have been refined taking into account the two described characteristics (Hu et al., 2023), i.e., the percentage of triples containing hyper-relational knowledge (mixed-percentage or fixed-percentage) and the number of qualifiers associated to main triples (mixed-qualifier or fixed-qualifier). We can thus construct three different sub-datasets, namely: Fixed-percentage Fixed-qualifier, Fixed-percentage Mixed-qualifier, and Mixed-percentage Mixed-qualifier. Note that the by definition Mixed-percentage Fixed-qualifier scenario does not exist. The statistics of the considered datasets are as shown in the Table 6.
  1. (1)
@@
  2. (2)
    Fixed-percentage Mixed-qualifier. Following (Galkin et al., 2020; Hu et al., 2023), we construct subsets where the number of qualifiers for each triple varies, but the proportion of triples with hyper-relational knowledge is fixed. For example, in the subsets WD50K (33), WD50K (66), and WD50K (100), the number in parentheses indicates the corresponding proportion of main triples with hyper-relational knowledge. We consider similar subsets for the WikiPeople and JF17K datasets.
  3. (3)
    Mixed-percentage Mixed-qualifier. This is the original setting introduced in (Galkin et al., 2020), used to test the performance under variable number of main triples with qualifiers and with different number of qualifiers associated to those triples.
Table 6. Statistics of datasets under three different scenarios. The values in parentheses indicate the percentage of triples in the corresponding dataset with hyper-relational facts. The hyphen -n indicates that each main triple in the corresponding dataset contains n qualifier pairs.
@@
WD50K 166435 23913 46159 47155 531
WikiPeople 294439 37715 37712 34825 178
JF17K 76379 - 24568 28645 501
B   Baselines
We compare HyperMono with twelve state-of-the-art baselines for HKGC, including Embedding-based Methods, Transformer-based with GNNs Methods, and Transformer-based without GNNs Methods: